OpenStudy (anonymous):

Solve: 2 cos^2 x - 3 cos x + 1 = 0 for 0< x<2pi -

OpenStudy (anonymous):

How would you rewrite this equation if b=cos(x)? and you subbed b for every instance of cos(x)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

\[2\cos(x)\cos(x) -3\cos(x) +1 =0\] so what does this equation become when you replace every instance of cos(x) with the letter b?

OpenStudy (anonymous):

Do you still need help with this question?

OpenStudy (anonymous):

Please reply so I would know If i should continue.

OpenStudy (anonymous):

you can continue

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

Okay, sorry I was helping someone else. So Start by identifying what this equation really is ...

OpenStudy (anonymous):

This is an ax^2+bx+c problem

OpenStudy (anonymous):

Which means that it might be possible to factor so.... 2 cos^2 x - 3 cos x + 1 = 0 ------> (cosx-1/2)(cosx-1)=0 is the factored form Let me know if I should continue.

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

Now, this also means that the factors can be rewritten as... (2cosx-1)(cosx-1)=0 Now solve the first equation (2cosx-1)=0------> 2cosx-1=0 Add 1, 2cosx=1 Divide by 2 cosx=1/2

OpenStudy (anonymous):

Now take the arccos(1/2). There is only solution from 0 to 2pi. Which is ---> pi/3 Now solve the second equation.

OpenStudy (anonymous):

Actually, sorry, there would be 2 solution for cosx=1/2 p/3, and 5pi/3 Now, solve the second equation.

OpenStudy (anonymous):

Do you follow so far?

OpenStudy (anonymous):

Continue?

OpenStudy (anonymous):

(cosx-1), now solve this for x, Add 1 cosx=1 Take the arccos(1) You get 2pi